OpenStudy (anonymous):

Select the pair of equations whose graphs are perpendicular. A. 2x – 8y = 9 12x – 3y = 7 B. 3x + 6y = 8 y = 2x – 8 C. y = 2x – 7 x + y = 3 D. 2y = –3x + 5 2x + 3y = 4

OpenStudy (anonymous):

I need a response like ASAP

OpenStudy (shamil98):

Perpendicular equations have inversed slopes. Change the the equations and compare which one has the inverse slopes. Example of perpendicular equations: y = 3x + 5 y = -1/3x + 5

OpenStudy (anonymous):

@shamil98 How can i change the equations though??

OpenStudy (shamil98):

Solve for y. This will get them into slope-intercept form. y = mx + b where m = your slope.

OpenStudy (anonymous):

okayy thank you